Explain the terms stress, strain, modulus of elasticity and elastic limit. Derive an expression in terms of tensile force and extension for the energy stored in a stretched rubber cord which obeys Hooke’s law.
1
Expert's answer
2020-10-29T07:03:10-0400

Tensile stress is a type of longitudinal deformation of a rod or rod that occurs when a load is applied to it along its longitudinal axis (the resultant of the forces acting on it is perpendicular to the cross section of the rod and passes through its center of mass).Measured in Pascals.

Elongation-deformation in which the load is applied longitudinally to the body, that is, coaxially or parallel to the points of attachment of the body, equal to the ratio between the change in length and the initial length. Measured in meters.

Young's modulus or tensile modulus is a mechanical property that measures the tensile stiffness of a solid material. They are expressed in Pascals.

Elastic limit — a property of a substance, the maximum load stress, after removal of which there are no residual (plastic) deformations.

Tensile force: "F=E\times S \times \varepsilon" , "E" is Young's modulus, "S" is cord cross-sectional area, "\varepsilon" is Elongation-deformation.

Cord energy: "E=\frac{k\times(\Delta l)^2}{2}=\frac{\frac{E\times S}{l}\times(\Delta l)^2}{2}" , "k" is elastic coefficient, "\Delta l" is changing cord length.